Let X represent the sum of the integers between 1 and 300, inclusive, and let Y represent the sum of the integers between 250 and 300, inclusive. Compute X-Y.
Answer by richard1234(7193)

X-Y = (1+2+...+300) - (250+251+...300) The 250, 251, ...300 all cancel, leaving

X-Y = 1+2+...+249 = 249*250/2 = 31125